Comparative financial statements for Weaver Company follow:

Weaver Company
Comparative Balance Sheet
at December 31
This Year Last Year
Assets
Cash $ 16 $ 22
Accounts receivable 490 320
Inventory 165 210
Prepaid expenses 10 8
Total current assets 681 560
Property, plant, and equipment 630 520
Less accumulated depreciation 120 70
Net property, plant, and equipment 510 450
Long-term investments 13 42
Total assets $ 1,204 $ 1,052
Liabilities and Stockholders' Equity
Accounts payable $ 340 $ 260
Accrued liabilities 40 50
Income taxes payable 79 72
Total current liabilities 459 382
Bonds payable 295 220
Total liabilities 754 602
Common stock 321 400
Retained earnings 129 50
Total stockholders’ equity 450 450
Total liabilities and stockholders' equity $ 1,204 $ 1,052
Weaver Company
Income Statement
For This Year Ended December 31
Sales $ 820
Cost of goods sold 460
Gross margin 360
Selling and administrative expenses 233
Net operating income 127
Nonoperating items:
Gain on sale of investments $ 9
Loss on sale of equipment (6 ) 3
Income before taxes 130
Income taxes 39
Net income $ 91

During this year, Weaver sold some equipment for $13 that had cost $43 and on which there was accumulated depreciation of $24. In addition, the company sold long-term investments for $38 that had cost $29 when purchased several years ago. Weaver paid a cash dividend this year and the company repurchased $79 of its own stock. This year Weaver did not retire any bonds.

Required:

1. Using the direct method, adjust the company’s income statement for this year to a cash basis.

2. Using the information obtained in (1) above, along with an analysis of the remaining balance sheet accounts, prepare a statement of cash flows for this year.
Answer #1
1)

Sales$     820.00
Increase in accounts receivable$   (170.00)$     650.00
Cost of goods sold$     460.00

Adjustments to a cash basis:



Decrease  in inventory$     (45.00)
Increase in accounts payable$     (80.00)$     335.00
Selling and administrative expenses$     233.00

Adjustments to a cash basis:



Increase in prepaid expenses

$         2.00
Decrease in accrued liabilities$       10.00
Depreciation charges$     (74.00)$     171.00
Income taxes$       39.00

Adjustments to a cash basis:



Increase in deferred income taxes$       (7.00)$       32.00
Net cash provided by operating activities$     112.00



2)

Weavers Company

Statement of Cash Flows

For the Year Ended December 31

Operating activities:

Cash received from customers
$     650.00
Less cash disbursements for:

Cost of merchandise purchased$     335.00
Selling and administrative expenses$     171.00
Income taxes$       32.00
Total cash disbursements
$     538.00
Net cash provided by operating activities
$     112.00
Investing activities:

Proceeds from sale of long-term investments38
Proceeds from sale of equipment13
Additions to plant and equipment-103
Net cash used for investing activities
-52
Financing activities:

Increase in bonds payable75
Decrease in common stock-79
Cash dividends-62
Net cash provided by financing activities
-66
Net decrease in cash
$       (6.00)
Cash balance, January 1
$       22.00
Cash balance, December 31
$       16.00


Last Year 490 165 22 320 210 560 510 520 70 450 42 1052 340 260 79 This Year 6 Assets 7 Cash 16 8 Accounts receivable 9 Inven
=C31 =C8-B8 =C32 =B47+B48 -45 =B49+B51+B52 =C18-B18 233 =B10-C10 =C19-B19 --(B13-C13+24) =C40 ESUM(B53:357) =C20-B20 =B58+B60


answered by: ANURANJAN SARSAM